# Explaining Grokking and Information Bottleneck through Neural Collapse Emergence

### v1

@@ 5.2 IB Dynamics
In this section, we conduct experiments on IB dynamics. Using the same setup as in the previous experiments, we measure how the MI between the learned representation $Z$ and the input $X$, as well as between $Z$ and the target $Y$, evolves during training. A fundamental difficulty in IB experiments with DNNs is that, beyond toy settings, conventional estimators based on binning or kernel density estimation fail to provide accurate estimates in high-dimensional input or representation due to the curse of dimensionality. This challenge remains an active research area, and in this work, we adopt the recent dimensionality-reduction-based MI estimator of Butakov et al. (2024b). Specifically, we first compress the variables into four dimensions and then estimate each entropy term of MI using the k-NN-based Kozachenko-Leonenko method (Kozachenko, 1987; Berrett et al., 2019). We denote this estimate by $\hat{I}$ and provide the details in Appendix B. In addition, to support the reliability of our experimental findings and to address the inherent difficulty of MI estimation, we use the normalized Hilbert-Schmidt independence criterion (nHSIC) (Gretton et al., 2005), a proxy widely adopted in information-theoretic analysis of DNNs. Please see Appendix B for the background.
Figure 4 shows, under the same setting as in Figure 3, the behavior of redundant information in the IB principle discussed in Section 3.2 together with the corresponding RNC1 score. As shown in the grokking experiments, the decrease of the RNC1 score in the later training stage occurs earlier when the weight decay is stronger (right), which is consistent with Theorem 4.3. Theorems 3.4 and 4.1 establish that this decrease in the RNC1 score contributes to the reduction of the redundant information, and the figure demonstrates this result. The left figure shows MI estimates, indicating that the stronger weight decay accelerates the decrease of redundant information. Although the information reduction is slightly delayed for large weight decay values ($\lambda=0.3,0.1$), the decrease of RNC1 scores actually leads to the reduction of redundant information. Since MI is estimated by decomposing it into differential entropy terms that are estimated separately, the resulting MI estimates can take negative values despite the non-negativity of MI, while still capturing the overall decreasing trend. To further support our findings, we also include the results using nHSIC to measure superfluous information (middle). This result exhibits the same qualitative behavior as the RNC1 score and corroborates our theoretical analysis of the IB dynamics. Additional results for other model architectures and datasets are provided in Section C.2.
@@ 6 Conclusion
In this work, we focus on grokking and IB dynamics as two representative late-phase phenomena of DNNs, whose mechanisms have been elusive. We show that both phenomena can be explained in terms of the population within-class variance of the learned representations, and more specifically, by the progression of neural collapse and its associated time scale. These theoretical findings are supported by our experiments. A natural next step is to extend the time-scale analysis of neural collapse to other architectures beyond MLP or to different initialization methods. Another interesting direction is to analyze the possibility of neural collapse that implicitly arises without weight decay.
Appendix A Proof of Main Results
A.1 Grokking Results
@@
Lemma A.1 (Cantelli’s inequality).
@@
where we used union bound in the last argument. From Lemma A.1, the term in the last line can be further bounded as follows:
@@
A.2 IB Dynamics
We first show the derivation of Proposition 3.3.
@@ A.2 IB Dynamics
Lemma A.2 (Cover & Thomas (2006), Theorem 8.6.5).
@@
which again follows from the Markov chain. For the second term, we use the differential entropy of Gaussian distribution (Cover & Thomas, 2006, Theorem 8.4.1), leading to
@@
Then, from Lemma A.2, the first term in Equation 17 can be computed as follows:
@@
A.3 Concentration of Within-Class Variance
@@
A.4 Neural Collapse Dynamics
A.4.1 Preliminaries
We first provide the formal assumptions used in Theorem 4.3, adopted from Jacot et al. (2025). While the validity of these assumptions is discussed in their work, we briefly comment on them here for completeness. Assumptions A.3 and A.4 are commonly made in the literature on gradient descent optimization, including (Nguyen & Hein, 2017; Nguyen & Mondelli, 2020; Frei et al., 2022). Assumption A.5 holds when the second layer is initialized with small values and the other parameters with sufficiently large values.
Assumption A.3 (Pyramidal network).
@@ A.4.1 Preliminaries
Assumption A.4 (Smooth activation).
@@
A.4.2 Proof of Theorem 4.3
@@
Proposition A.6 (Theorem B.1 of Jacot et al. (2025)).
@@
Corollary A.7.
Under the conditions of Proposition A.6, if $\xi_{1}\leq\frac{1}{2}\min_{c\in[K]}\sqrt{n_{c}}$, we have
@@
Proof of Corollary A.7.
By the definition of RNC1, it suffices to bound $\operatorname{Tr}({\bm{\Sigma}}_{W})$ from above and $B_{g}$ from below. Substituting $\sigma_{K}({\bm{Y}})=\min_{c\in[K]}\sqrt{n_{c}}$, which follows from ${\bm{Y}}{\bm{Y}}^{\top}=\operatorname{diag}\left(n_{1},\ldots,n_{K}\right)$, into Proposition A.6, we have
@@
Combining Equations 35 and 36 yields the desired upper bound. Finally, regarding the condition on $\xi_{1}$, we used that $\sigma_{K}({\bm{Y}})=\min_{c\in[K]}\sqrt{n_{c}}\leq\sqrt{N/K}$ and $K\geq 2$.
Proposition A.8.
Suppose that the network $f$ satisfies Assumptions A.3, A.4 and A.5 and that the input domain $\mathcal{X}$} is bounded, i.e., $\|{\bm{x}}\|_{2}\leq B_{x}$ for all ${\bm{x}}\in{\mathcal{X}}$. Fix $0<\epsilon_{1}<\frac{1}{2\sqrt{2}}\min_{c\in[K]}\sqrt{n_{c}}$ and $\epsilon_{2}>0$. Let weight decay parameter $\lambda$ and step size $\eta$ satisfy the following:
@@
Proof of Proposition A.8.
The proof follows the same argument as in the proof of Jacot et al. (2025, Theorem B.2). Under Assumptions A.3, A.4 and A.5, the PL property of the unregularized loss $\widehat{\mathcal{L}}_{0}({\bm{\theta}})$ is established. While the presence of a weight-decay term slightly shifts the convergence point, the PL condition still holds for the regularized loss $\widehat{\mathcal{L}}_{\lambda}({\bm{\theta}})$. From the PL condition of $\widehat{\mathcal{L}}_{\lambda}({\bm{\theta}})$ around the initialization, there exists a time step
@@
and for any time step $\tau\geq\tau_{2}$, we have $\|{\bm{W}}_{L}(\tau)^{\top}{\bm{W}}_{L}(\tau)-{\bm{W}}_{L-1}(\tau){\bm{W}}_{L-1}(\tau)^{\top}\|_{2}\leq\epsilon_{2}$. Applying Corollary A.7 with $\xi_{1}=\sqrt{2}\epsilon_{1}$, $\xi_{2}=\epsilon_{2}$, and the upper bound $r$ yields the desired result. ∎
@@
The conclusion follows from Proposition A.8; specifically, by replacing $\epsilon_{1}^{2}$ with $\epsilon_{1}$ and evaluating the order of $\tau_{1}^{\prime}$ and $\tau_{2}$ in the proposition. By the condition of $\lambda\leq\alpha$ in Proposition A.8 and the definition of $m_{\lambda}$, we can bound $m_{\lambda}$ as a constant order depending only on the initialization. Consequently, the two terms in the upper bound of RNC1 are each multiplied by scales independent of $\lambda$ and $\mu$; thus, we have $\mathrm{RNC1}(\tau)=O\left((\sqrt{\epsilon_{1}}+\sqrt{\epsilon_{2}})^{2}\right)=O\left(\epsilon_{1}+\epsilon_{2}\right)$. Finally, we evaluate the order of the upper bounds on $\tau_{1}^{\prime}$ and $\tau_{2}$. Since we have $\log(1-x)\approx-x$ for small $x$, the desired order expression of time steps is obtained.
Appendix B Experimental Details
@@
Even approaches based on autoencoders require assumptions such as invertibility for accurate estimation and demand additional computational resources. To further support the findings of our study, we also conducted experiments based on normalized HSIC. Fukumizu et al. (2007, Theorem 4) shows that a normalized version of HSIC (to be defined below) coincides with the chi-square divergence between the joint distribution $P_{X,Y}$ and the product of marginals $P_{X}P_{Y}$. Given that the MI is defined using the KL divergence instead of chi-square divergence, this suggests that the HSIC-based quantity can be interpreted as a variant of MI. More specifically, Gibbs & Su (2002, Theorem 5) shows that the chi-square divergence upper-bounds the KL divergence. Using HSIC as an information-theoretic quantity is a common practice in the literature on IB and learning algorithms (Ma et al., 2020; Pogodin & Latham, 2020; Wang et al., 2021; Jian et al., 2022; Guo et al., 2023; Wang et al., 2023; Sakamoto & Sato, 2024).
@@
Appendix C Additional Experiments
C.1 Grokking

### v2

@@ 5.2 IB Dynamics
In this section, we conduct experiments on IB dynamics. Using the same setup as in the previous experiments, we measure how the MI between the learned representation $Z$ and the input $X$, as well as between $Z$ and the target $Y$, evolves during training. A fundamental difficulty in IB experiments with DNNs is that, beyond toy settings, conventional estimators based on binning or kernel density estimation fail to provide accurate estimates in high-dimensional input or representation due to the curse of dimensionality. This challenge remains an active research area, and in this work, we adopt the recent dimensionality-reduction-based MI estimator of Butakov et al. (2024b). Specifically, we first compress the variables into four dimensions and then estimate each entropy term of MI using the k-NN-based Kozachenko-Leonenko method (Kozachenko, 1987; Berrett et al., 2019). We denote this estimate by $\hat{I}$ and provide the details in Appendix C. In addition, to support the reliability of our experimental findings and to address the inherent difficulty of MI estimation, we use the normalized Hilbert-Schmidt independence criterion (nHSIC) (Gretton et al., 2005), a proxy widely adopted in information-theoretic analysis of DNNs. Please see Appendix C for the background.
Figure 4 shows, under the same setting as in Figure 3, the behavior of redundant information in the IB principle discussed in Section 3.2 together with the corresponding RNC1 score. As shown in the grokking experiments, the decrease of the RNC1 score in the later training stage occurs earlier when the weight decay is stronger (right), which is consistent with Theorem 4.3. Theorems 3.4 and 4.1 establish that this decrease in the RNC1 score contributes to the reduction of the redundant information, and the figure demonstrates this result. The left figure shows MI estimates, indicating that the stronger weight decay accelerates the decrease of redundant information. Although the information reduction is slightly delayed for large weight decay values ($\lambda=0.3,0.1$), the decrease of RNC1 scores actually leads to the reduction of redundant information. Since MI is estimated by decomposing it into differential entropy terms that are estimated separately, the resulting MI estimates can take negative values despite the non-negativity of MI, while still capturing the overall decreasing trend. To further support our findings, we also include the results using nHSIC to measure superfluous information (middle). This result exhibits the same qualitative behavior as the RNC1 score and corroborates our theoretical analysis of the IB dynamics. Additional results for other model architectures and datasets are provided in Section D.2.
@@ 6 Conclusion
In this work, we focus on grokking and IB dynamics as two representative late-phase phenomena of DNNs, whose mechanisms have been elusive. We show that both phenomena can be explained in terms of the population within-class variance of the learned representations, and more specifically, by the progression of neural collapse and its associated time scale. These theoretical findings are supported by our experiments. Beyond the theoretical perspective, our results also provide practical implications: tracking quantities such as the rescaled within-class variance can help determine when continued training will be beneficial, and weight decay can accelerate the transition to this late-phase regime. A natural next step is to extend the time-scale analysis of neural collapse to other architectures beyond MLP or to different initialization methods. Another interesting direction is to analyze the possibility of neural collapse that implicitly arises without weight decay.
7 Acknowledgments
This work was supported by JSPS KAKENHI Grant Number JP24H00709. We thank the reviewers for their constructive comments and the members of our laboratory for helpful discussions.
Appendix A Further Comparison with Related Work
As a concurrent work to ours, we compare our results with Han et al. (2025) in this section. Since the title may appear to contradict our claims at first glance, we clarify that their findings are fully consistent with ours and highlight the novelty of our contribution. Han et al. (2025) discusses grokking in relation to neural collapse, which makes their setting similar to ours. Although the title suggests that neural collapse might not be relevant for generalization, this is not what the paper actually argues. Rather, they state that neural collapse appears earlier than the acquisition of generalization. The emergence of neural collapse does not align with that of generalization, and therefore it does not fully explain generalization ability. In contrast, they observe that flatness correlates more closely with the decrease in test loss and thus appears to offer a better explanation in their experiments. However, this empirical observation is entirely consistent with our analysis. The apparent contradiction arises because Han et al. (2025) measures neural collapse using the neural collapse clustering (NCC) metric, which mixes multiple properties associated with neural collapse. In particular, NCC also captures the separation of class mean representations that naturally occurs through fitting the training set, and therefore it decreases earlier than the test loss. As we emphasize in Remark 4.2, our analysis isolates the contribution of within-class variance or RNC1 score, allowing us to separate out the cause of the generalization and correctly focus on the variance decrease that is actually responsible for the emergence of generalization in grokking. Furthermore, the theoretical analysis in Han et al. (2025) is limited to showing that flatness is partially guaranteed under neural collapse. Their work addresses neither the connection to generalization nor the associated training dynamics. In contrast, as illustrated in Figure 1, our paper clarifies the relationships among several concepts, including information bottleneck, and it further characterizes their training dynamics through a neural-collapse-based analysis. This leads to a unified understanding of the mechanisms underlying grokking.
Appendix B Proof of Main Results
B.1 Grokking Results
@@
Lemma B.1 (Cantelli’s inequality).
@@
where we used union bound in the last argument. From Lemma B.1, the term in the last line can be further bounded as follows:
@@
B.2 IB Dynamics
B.2.1 Proof of Proposition 3.3
@@ B.2 IB Dynamics
Lemma B.2 (Cover and Thomas (2006), Theorem 8.6.5).
@@
which again follows from the Markov chain. For the second term, we use the differential entropy of Gaussian distribution (Cover and Thomas, 2006, Theorem 8.4.1), leading to
@@
Then, from Lemma B.2, the first term in Equation 17 can be computed as follows:
@@
B.2.3 Remark on Upper-bound Tightness
In Theorem 3.4, the redundant information $I(Z;X\mid Y)$ is upper-bounded with a population within-class variance. In this section, we discuss how informative this upper bound is.
From the proof of Theorem 3.4, we obtain the following tight bound on $I(Z;X\mid Y)$:
$I(Z;X\mid Y)$ $\leq\frac{1}{2}\mathbb{E}_{y\sim Y}\left[\log\det\left(\frac{\mathrm{Cov}_{X\mid Y=y}[\tilde{g}(X)]}{\sigma^{2}}+{\bm{I}}\right)\right].$ (26)
This bound is derived from Lemma B.2, and equality holds when $Z\mid Y=y$ follows the Gaussian distribution. Therefore, this upper bound is tight. In the subsequent step of the proof, we used the inequality $\log\det({\bm{A}}+{\bm{I}})\leq\operatorname{Tr}({\bm{A}})$. Since the equality does not hold when $\bm{A}$} is positive definite, the resulting bound in Theorem 3.4 is not tight. However, regarding the within-class variance term appearing in the theorem, we can show that reducing the variance in each coordinate always decreases both the tight bound above and the upper bound in Theorem 3.4
Proposition B.3.
Reducing the population within-class variance in any single coordinate, i.e., $\mathbb{E}_{X\mid Y=y}\left[\left(\tilde{g}(X)-\mathbb{E}_{X\mid Y=y}[\tilde{g}(X)]\right)_{i}^{2}\right]$ for any $i\in[d_{rep}]$ and $y\in[K]$, strictly decreases both the upper bound in Theorem 3.4 and the tight upper bound on $I(Z;X\mid Y)$ in Equation 26.
Proof of Proposition B.3.
For the upper bound in Theorem 3.4, since it can be written as
$\sum_{i\in[d_{rep}]}\mathbb{E}_{Y}\mathbb{E}_{X\mid Y}\left[\left(\tilde{g}(X)-\mathbb{E}_{X\mid Y}[\tilde{g}(X)]\right)_{i}^{2}\right],$ (27)
it is obvious that decreasing any summand decreases the upper bound. We now show that the tight upper bound in Equation 26 also decreases under such a perturbation. Let the amount of the variance reduction along coordinate $i$ be $\delta>0$. For notational simplicity, we denote the original covariance matrix by ${\bm{K}}=\mathrm{Cov}_{X\mid Y}[\tilde{g}(X)]$ and the perturbed covariance by ${\bm{K}}^{\prime}={\bm{K}}-\delta{\bm{e}}_{i}{\bm{e}}_{i}^{\top}$. Then, the value of the upper bound is bounded below as
$\frac{1}{2}\mathbb{E}_{Y}\left[\log\det\left(\frac{{\bm{K}}^{\prime}+\delta{\bm{e}}_{i}{\bm{e}}_{i}^{\top}}{\sigma^{2}}+{\bm{I}}\right)\right]$ (28)
$=\frac{1}{2}\mathbb{E}_{Y}\left[\log\left\{\left(1+\frac{\delta}{\sigma^{2}}{\bm{e}}_{i}^{\top}\left(\frac{{\bm{K}}^{\prime}}{\sigma^{2}}+{\bm{I}}\right)^{-1}{\bm{e}}_{i}\right)\det\left(\frac{{\bm{K}}^{\prime}}{\sigma^{2}}+{\bm{I}}\right)\right\}\right]$ (29)
$>\frac{1}{2}\mathbb{E}_{Y}\left[\log\det\left(\frac{{\bm{K}}^{\prime}}{\sigma^{2}}+{\bm{I}}\right)\right],$ (30)
where the equality holds from the matrix determinant lemma, and the inequality follows from the fact that the diagonal element of the inverse of a positive definite matrix is always positive. Since the right-hand side is exactly the upper bound in Equation 26 evaluated at the perturbed covariance, this completes the proof. ∎
B.3 Concentration of Within-Class Variance
@@
B.4 Neural Collapse Dynamics
B.4.1 Preliminaries
We first provide the formal assumptions used in Theorem 4.3, adopted from Jacot et al. (2025). A detailed discussion of their validity and typicality is given in Remark B.6.
Assumption B.3 (Pyramidal network).
@@ B.4.1 Preliminaries
Assumption B.4 (Smooth activation).
@@
Remark B.6 (Validity of the Assumptions).
The derivation of neural collapse from gradient descent, rather than from the unconstrained feature model (UFM), was first established in Jacot et al. (2025), and Theorem 4.3 adopts the same assumptions. To analyze the convergence of the DNN training loss, one typically imposes conditions ensuring a positive lower bound on the Jacobian with respect to the parameters. There are multiple well-established ways to guarantee this, including width requirements across all layers (Du et al., 2019; Allen-Zhu et al., 2019; Zou and Gu, 2019) and other pyramidal-topology-based conditions with mild width assumptions (Nguyen and Mondelli, 2020; Bombari et al., 2022; Karhadkar et al., 2024). Our Assumptions B.3, B.4 and B.5 can be replaced by any of these alternatives. For example, Allen-Zhu et al. (2019); Zou and Gu (2019) analyze ReLU networks instead of smooth activations but require all layers to be sufficiently wide. In contrast, the pyramidal topology assumption (Assumption B.3) removes the need for width assumptions on every layer and replaces them with a more realistic setting: a wide first layer followed by a narrowing architecture. The smooth-activation assumption (Assumption B.4) is widely used in convergence analysis and appears in independent work such as Nguyen and Mondelli (2020); Bombari et al. (2022); Liu et al. (2022a); Frei et al. (2022). Smooth leaky ReLU satisfies this assumption and can approximate ReLU arbitrarily well for suitable choices of $\gamma$ and $\beta$. Assumption B.5 concerns initialization, and it can be satisfied by choosing the second-layer scale sufficiently small. Moreover, by relaxing Assumption B.3 from $d_{1}\geq N$ to $d_{1}=\Omega(N)$, Assumption B.5 can be shown to hold under standard He/LeCun initialization (LeCun et al., 2002; He et al., 2015), which follows directly from Appendix C of Nguyen and Mondelli (2020). Finally, the use of squared loss is standard in this line of work: it is used both in the gradient-descent NTK-style analyses (Du et al., 2019; Allen-Zhu et al., 2019; Zou and Gu, 2019; Nguyen and Mondelli, 2020; Jacot et al., 2025) and in theoretical neural collapse literature (Han et al., 2022; Zhou et al., 2022a; Súkeník et al., 2023; 2024). As noted in Jacot et al. (2025), once the training set is interpolated, the second part of the analysis (the time step $\tau_{2}$ in Theorem 4.3) showing the emergence of neural collapse does not rely on these assumptions. We note that none of our results except Theorem 4.3 depend on Assumptions B.3, B.4 and B.5.
B.4.2 Proof of Theorem 4.3
@@
Proposition B.7 (Theorem B.1 of Jacot et al. (2025)).
@@
Corollary B.8.
Under the conditions of Proposition B.7, if $\xi_{1}\leq\frac{1}{2}\min_{c\in[K]}\sqrt{n_{c}}$, we have
@@
Proof of Corollary B.8.
By the definition of RNC1, it suffices to bound $\operatorname{Tr}({\bm{\Sigma}}_{W})$ from above and $B_{g}$ from below. Substituting $\sigma_{K}({\bm{Y}})=\min_{c\in[K]}\sqrt{n_{c}}$, which follows from ${\bm{Y}}{\bm{Y}}^{\top}=\operatorname{diag}\left(n_{1},\ldots,n_{K}\right)$, into Proposition B.7, we have
@@
Combining Equations 40 and 41 yields the desired upper bound. Finally, regarding the condition on $\xi_{1}$, we used that $\sigma_{K}({\bm{Y}})=\min_{c\in[K]}\sqrt{n_{c}}\leq\sqrt{N/K}$ and $K\geq 2$.
Proposition B.9.
Suppose that the network $f$ satisfies Assumptions B.3, B.4 and B.5 and that the input domain $\mathcal{X}$} is bounded, i.e., $\|{\bm{x}}\|_{2}\leq B_{x}$ for all ${\bm{x}}\in{\mathcal{X}}$. Fix $0<\epsilon_{1}<\frac{1}{2\sqrt{2}}\min_{c\in[K]}\sqrt{n_{c}}$ and $\epsilon_{2}>0$. Let weight decay parameter $\lambda$ and step size $\eta$ satisfy the following:
@@
Proof of Proposition B.9.
The proof follows the same argument as in the proof of Jacot et al. (2025, Theorem B.2). Under Assumptions B.3, B.4 and B.5, the PL property of the unregularized loss $\widehat{\mathcal{L}}_{0}({\bm{\theta}})$ is established. While the presence of a weight-decay term slightly shifts the convergence point, the PL condition still holds for the regularized loss $\widehat{\mathcal{L}}_{\lambda}({\bm{\theta}})$. From the PL condition of $\widehat{\mathcal{L}}_{\lambda}({\bm{\theta}})$ around the initialization, there exists a time step
@@
and for any time step $\tau\geq\tau_{2}$, we have $\|{\bm{W}}_{L}(\tau)^{\top}{\bm{W}}_{L}(\tau)-{\bm{W}}_{L-1}(\tau){\bm{W}}_{L-1}(\tau)^{\top}\|_{2}\leq\epsilon_{2}$. Applying Corollary B.8 with $\xi_{1}=\sqrt{2}\epsilon_{1}$, $\xi_{2}=\epsilon_{2}$, and the upper bound $r$ yields the desired result. ∎
@@
The conclusion follows from Proposition B.9; specifically, by replacing $\epsilon_{1}^{2}$ with $\epsilon_{1}$ and evaluating the order of $\tau_{1}^{\prime}$ and $\tau_{2}$ in the proposition. By the condition of $\lambda\leq\alpha$ in Proposition B.9 and the definition of $m_{\lambda}$, we can bound $m_{\lambda}$ as a constant order depending only on the initialization. Consequently, the two terms in the upper bound of RNC1 are each multiplied by scales independent of $\lambda$ and $\mu$; thus, we have $\mathrm{RNC1}(\tau)=O\left((\sqrt{\epsilon_{1}}+\sqrt{\epsilon_{2}})^{2}\right)=O\left(\epsilon_{1}+\epsilon_{2}\right)$. Finally, we evaluate the order of the upper bounds on $\tau_{1}^{\prime}$ and $\tau_{2}$. Since we have $\log(1-x)\approx-x$ for small $x$, the desired order expression of time steps is obtained. ∎
Appendix C Experimental Details
@@
Even approaches based on autoencoders require assumptions such as invertibility for accurate estimation and demand additional computational resources. To further support the findings of our study, we also conducted experiments based on normalized HSIC. Fukumizu et al. (2007, Theorem 4) shows that a normalized version of HSIC (to be defined below) coincides with the chi-square divergence between the joint distribution $P_{X,Y}$ and the product of marginals $P_{X}P_{Y}$. Given that the MI is defined using the KL divergence instead of chi-square divergence, this suggests that the HSIC-based quantity can be interpreted as a variant of MI. More specifically, Gibbs and Su (2002, Theorem 5) shows that the chi-square divergence upper-bounds the KL divergence. Using HSIC as an information-theoretic quantity is a common practice in the literature on IB and learning algorithms (Ma et al., 2020; Pogodin and Latham, 2020; Wang et al., 2021; Jian et al., 2022; Guo et al., 2023; Wang et al., 2023; Sakamoto and Sato, 2024).
@@
Appendix D Additional Experiments
D.1 Grokking
Decision Boundaries for Additional Class Pairs.
(a) (a.1) Overfitting phase (time step $\tau_{1}=16{,}000$): Class 4 vs. 5.
(b) (a.2) Convergence phase (time step $\tau_{2}=100{,}000$): Class 4 vs. 5.
(c) (b.1) Overfitting phase (time step $\tau_{1}=16{,}000$). Class 8 vs. 9.
(d) (b.2) Convergence phase (time step $\tau_{2}=100{,}000$): Class 8 vs. 9.
Figure 5: Margins of individual examples at two time steps during grokking. The model is a 4-layer MLP on the MNIST dataset. Here, we provide the corresponding plots for several class pairs other than the 0-1 pair shown in Figure 2.
Figure 2 showed how representations evolve during training on MNIST, but due to space limitations in the main text, we presented only the decision boundary between class 0 and $1$. To demonstrate that the same phenomenon occurs for other class pairs as well, Figure 5 shows results for two additional pairs: class $4$ vs. $5$ and class $8$ vs. $9$. For these pairs, we observe the same trend: during the overfitting phase, training examples are already almost separable but still exhibit large within-class variance. As training proceeds, the training examples become more separated and more tightly clustered. Consequently, test examples are better separated and more concentrated in representation space, leading to improved generalization.